a gas is filled in a cylinder.its temperature is increased by 20% and volume decreases byby 20%.if pressure remains constant, what percentage of mass of the gas leaks out ?

Grade:

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Ankit

PV=mRT

take log

logP +logV =logm +logR +logT

differentiate

δP/P  +δV/V =δm/m +δT/T

or

δP/P *100   +δV/V *100 =δm/m*100 +δT/T *100

here

δP =0

δV/V *100 =-20

δT/T *100 =20

put in above equation

δm/m*100 =-40

percentage leakage of mass is 40


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin